OpenStudy (anonymous):

Katie has a pair of earrings in the shape of congruent triangles. The measure of angle ABC is equal to the measure of angle DEF. What is the value of x?

OpenStudy (anonymous):

A 1 B 2 C 3 D 4

OpenStudy (hoblos):

2x-3 = x-1 x = 2

OpenStudy (anonymous):

x would be C, 3

OpenStudy (anonymous):

Lol i got B from one person and C from another lol

OpenStudy (hoblos):

@emecham ... how did you get the 3?!!

OpenStudy (anonymous):

The sides can't be the same or more then the hypotenuse right ? so it can't be 3 lol

OpenStudy (anonymous):

oh wait, my bad.... @hoblos is right.... sorry i misread it

OpenStudy (anonymous):

haha its fine !(: thanks tho !